Barbecue Industries produces two products, Barbecue sauce and Barbecue seasoning. Information relating to the products for OcPlease check my answers for tasks 2, 3, and 4, and add wording as needed. If you can double check for me the answers for number one would be great also. Thank you.
Answer #1
  1. The no.1 is correctly solved in all aspects
  2. If Barbeque sauce is dropped, then, only Barbeque seasoning will be produced.

The reported segment profit of Barbeque seasoning in requirement 1 is $25,000. The entire common expenses will be adjusted to this only.

The net income(loss) = 25,000 – 120,000 = $ (95,000)

The net loss will increase by 95,000 – 35,000 = $ 60,000

  1. If Barbeque seasoning is dropped, then, only Barbeque sauce will be produced.

The reported segment profit of Barbeque sauce in requirement 1 is $60,000. The entire common expenses will be adjusted to this only.

The net income(loss) = 60,000 – 120,000 = $ (60,000)

The net loss will increase by 60,000 – 35,000 = $ 25,000

  1. Barbeque seasoning should be dropped as the net loss will only decrease by $ 25,000 compared to $ 60,000 when Barbeque sauce is dropped

kindly upvote